graph the equation and identify the y intercept
y= 4/3X +2

Answer by mananth(12269)

graph the equation and identify the y intercept
y= 4/3X +2
...
x=0,y=2 (0,2)
y=0,x=-3/2(-3/2,0)
...

y intercept =2